what is the answer to U-12=0
Found 2 solutions by Alan3354, pragya.ban@gmail.com:
Answer by Alan3354(69443)

what is the answer to U-12=0
---------------
Depends on the question.
---
If you want to solve U-12 = 0 for U:
Add 12 to both sides.

Answer by pragya.ban@gmail.com(32)

what is the answer to U-12=0
U=12

Hope this helps! If not, feel free to ask again.